# Precision Determination of Scintillation Screen Parameters from Annual Modulation Measurement of Pulsar Scintillation Arc Curvature with the FAST Telescope

### v1

@@ 4.1 PSR B1237+25
Scintillation arcs B and E were only identified in observations on November 21, 2019, and January 18, 2020. The two dates correspond to days of year that are too close together, so we essentially have no annual modulation information for them. We applied the assumptions that $\text{V}_{{\text{screen},~{}\zeta}}$ is 0 and that $\mathbf{n_{\text{screen}}}$ is aligned with $\mathbf{V_{\text{eff}}}$ to estimate their fractional distance s. In this case, s will have two solutions: a solution close to 0, which is the upper limit when the screen is close to the pulsar, and a solution close to 1, which is the lower limit when the screen is close to Earth. For arc C, we only have two observed values that are close together, giving us little annual modulation information, so it was processed in the same way.
4.2 PSR B1842+14
@@
We conducted 4 observations of pulsar B1842+14, each identifying 1 scintillation arc, as shown in Table 4. The reason for observing only one arc may be its relatively low brightness, which is an order of magnitude lower than the other two pulsars studied in this paper (as shown in Table 1). Yao et al. (2020) observed 2 scintillation arcs from this pulsar on August 28, 2018, one with a curvature of 0.12 $\pm$ 0.05 s3 at 1400 MHz, which is 785 $\pm$ 327 $\mathrm{m}^{-1}\,\mathrm{mHz}^{-2}$. We compared this value with our results (as shown in Figure 9). Yao et al. (2020) also observed another scintillation arc with a curvature of 0.0010 $\pm$ 0.0003 s3 at 1400 MHz, which we did not observe. Possible reasons include that this arc has moved away from the line of sight to the pulsar, or it may be due to significant noise interference. Further research on this issue will require longer continuous observations to obtain data with higher signal-to-noise ratios. The scintillation arc we observed was relatively diffuse (as shown in Figure 4), resulting in larger errors in the curvature measurements. The limited dataset of 4 observations reduces the precision of our parameter-fitting results. We hope to conduct more observations of pulsar B1842+14 in the future to improve the modeling accuracy. We modeled the annual modulation of the scintillation arc (as shown in Figure 9). Through modeling the annual modulation, we determined that the fractional distance s of the scintillation screen is $0.85^{+0.05}_{-0.10}$, corresponding to a distance of $0.24^{+0.21}_{-0.12}$ kpc; the scattering direction $\zeta$ is $122^{+11}_{-12}$ ; and the velocity component along the scattering direction $\text{v}_{\text{scr}}$ is $19^{+11}_{-11}$ km/s. Yao et al. (2020) reported a result of 0.3 $\pm$ 0.2 kpc for this arc based on a single observation assuming $\mathbf{n_{\text{screen}}}$ aligns with $\mathbf{V}_{\text{eff}}$ and $\text{V}_{\text{screen}}$ = 0. The large uncertainty for these screen distance measurements is mainly attributed to the uncertainty in the pulsar distance. Accurate pulsar distance measurements are vital for obtaining more precise screen distance determinations.
4.3 PSR B2021+51
@@
We discovered at least 6 scintillation arcs in the observation data of pulsar B2021+51 (see Table 5), with 6 scintillation arcs identifiable in the observation data from September 17, 2022, as shown in Figure 5. In contrast, Stinebring et al. (2022) observed only 2 scintillation arcs when studying the scintillation phenomenon of this pulsar using the Green Bank Telescope. This demonstrates that the FAST telescope possesses exceptional observational capabilities and is highly suitable for studying pulsar scintillation.
@@
Among these, the scintillation arc E was the brightest. Our model fitting yielded the following three parameters for this scintillation screen: a fractional distance s of $0.556^{+0.041}_{-0.038}$ (corresponding to a distance of $887^{+167}_{-132}$ pc), a scattering direction of $20.3^{+4.3}_{-4.3}$ , and a projected velocity of $-12.1^{+3.7}_{-3.8}$ km/s. For scintillation arc D, we also successfully fitted all three parameters of the scintillation screen.
@@
4.4 Spatial distribution of the scintillation screens
@@
Screen A and one of the alternative solutions for screens B and C in the direction of PSR B1237+25 are close to PSR B1237+25. Screens in proximity to pulsars may be associated with pulsar-local environments, such as the pulsar wind nebula and bow shocks. However, the characteristic age ($\dot{\text{P}}$) of PSR B1237+25 is 21 Myr (Mamatali et al., 2021), indicating that there is no pulsar wind nebula around this pulsar. If other structures, for example, a bow shock exists around PSR B1237+25, it remains challenging to associate these screens with such a structure. The screen most likely to be closest to PSR B1237+25 is screen A, with an estimated distance of $37.6^{+3.9}_{-3.4}$ pc from the pulsar (s is $0.0435_{-0.0027}^{+0.0031}$). This value was calculated based on two assumptions ($\text{V}_{{\text{screen},~{}\zeta}}$ is 0 and $\mathbf{n_{\text{screen}}}$ is aligned with $\mathbf{V_{\text{eff}}}$), thus representing an upper limit on the distance from the pulsar. Consequently, we only know that screen A is located $\lesssim$ 40 pc of the pulsar, which provides insufficient constraint to confidently associate it with potential pulsar-local structures. Similarly, PSR B2021+51 has a characteristic age of 2.75 Myr (Vlemmings et al., 2004) and lacks a pulsar wind nebula. The screen potentially closest to PSR B2021+51 is estimated to be located $\lesssim$ 70 pc, which also provides inadequate positional precision to establish a definitive connection with potential pulsar-local structures.
@@
Only three potential screens (alternative solutions for B, C, and E of PSR B1237+25) might reside within the Local Bubble, and these represent just one of two possible solutions for each screen, with the alternative placements being at much greater distances. These extremely nearby solutions, if physical, would be associated with very local interstellar structures at distances of only $<$ 1 pc. In general, scintillation screens within the Local Bubble typically produce very large curvature values, requiring extended observation times to properly resolve. Our moderate observation durations may have limited our ability to detect additional local screens.
Screens E and D for PSR 2021+51 are clearly unrelated to either the Local Bubble structure or the pulsar local environments, suggesting they originate from general interstellar medium structure distributed along the line of sight.
Because the pulsars in our sample are relatively distant, most of the screens we detected are also at considerable distances. The distribution of screens we observe is thus biased toward more distant structures, with potentially many closer screens remaining undetected due to the challenges in resolving their very high curvature values with our current observation strategies.
5 Conclusion
@@
Our findings reveal that these scintillation phenomena occur throughout various regions of interstellar space between Earth and the observed pulsars. This distribution pattern indicates that electron density fluctuations capable of causing such scintillation are widespread throughout interstellar space. For screens located in close proximity to pulsars, our current positional constraints are insufficient to confidently associate them with pulsar-local environments. Several screens in our sample appear at distances that coincide with the Local Bubble boundary, particularly the brightest scintillation arc in the direction of PSR B1237+25. When confirmed to be associated with the Local Bubble boundary, the distance measurement of such scintillation screens would provide a valuable constraint on the Local Bubble shape at high Galactic latitudes.
@@
6 Acknowledgments
This work made use of the data from FAST (Five-hundred-meter Aperture Spherical radio Telescope). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences. This work is supported by the National Natural Science Foundation of China grant No. 12373025. Jumei Yao was supported by the National Science Foundation of Xinjiang Uygur Autonomous Region (2022D01D85), the Major Science and Technology Program of Xinjiang Uygur Autonomous Region (2022A03013-2), the Tianchi Talent Project, and the CAS Project for Young Scientists in Basic Research (YSBR-063), the Tianshan Talents Program (2023TSYCTD0013), and the Chinese Academy of Sciences (CAS) “Light of West China” Program (No. xbzg-zdsys-202410 and No. 2022-XBQNXZ-015). Weiwei Zhu is supported by the National SKA Program of China No. 2020SKA0120200, the National Natural Science Foundation grant No. 12041303. Yonghua Xu is supported by the National Key R&D Program of China (grant No. 2022YFC2205203), the CAS “Light of West China” Program. Xun Shi acknowledges helpful discussions with Kejia Lee and Yulan Liu.
@@ 6 Acknowledgments
Appendix A Appendix information
@@
Figure 11: The posterior distributions of parameters for screen G in the direction of PSR B1237+25, as well as the parallax and proper motion of PSR B1237+25. In these figures, blue solid lines indicate the median, vertical dashed lines indicate 1 $\sigma$ confidence levels, and contours show 1 $\sigma$, 2 $\sigma$, and 3 $\sigma$ confidence levels.
Figure 12: The posterior distributions of parameters for the screen in the direction of PSR B1842+14, as well as the distance and proper motion of PSR B1842+14. In these figures, blue solid lines indicate the median, vertical dashed lines indicate 1 $\sigma$ confidence levels, and contours show 1 $\sigma$, 2 $\sigma$, and 3 $\sigma$ confidence levels.
Figure 13: The posterior distributions of parameters for screen E in the direction of PSR B2021+51, as well as the parallax and proper motion of PSR B2021+51. In these figures, blue solid lines indicate the median, vertical dashed lines indicate 1 $\sigma$ confidence levels, and contours show 1 $\sigma$, 2 $\sigma$, and 3 $\sigma$ confidence levels.

### v2

@@ IV.1 PSR B1237+25
Scintillation arcs B and E were only identified in observations on November 21, 2019, and January 17, 2020. The two dates correspond to days of the year that are too close together, so we essentially have no annual modulation information for them. We applied the assumptions that $v_{{\text{screen},~\zeta}}$ is 0 and that $\mathbf{n}_{\text{screen}}$ is aligned with $\mathbf{v}_{\text{eff}}$ to estimate their fractional distance $s$. In this case, $s$ will have two solutions: a solution close to 0, which is the upper limit when the screen is close to the pulsar, and a solution close to 1, which is the lower limit when the screen is close to Earth. For arc C, we only have two observed values that are close together, giving us little annual modulation information, so it was processed in the same way.
IV.2 PSR B1842+14
@@
We conducted 4 observations of pulsar B1842+14, each identifying 1 scintillation arc, as shown in Table 4. The reason for observing only one arc may be its relatively low brightness, which is an order of magnitude lower than the other two pulsars studied in this paper (as shown in Table 1). Yao et al. (2020) observed 2 scintillation arcs from this pulsar on August 28, 2018, one with a curvature of 0.12 $\pm$ 0.05 s3 at 1400 MHz, which is 785 $\pm$ 327 $\mathrm{m}^{-1}\,\mathrm{mHz}^{-2}$. We compared this value with our results (as shown in Figure 10). Yao et al. (2020) also observed another scintillation arc with a curvature of 0.0010 $\pm$ 0.0003 s3 at 1400 MHz, which we did not observe. Possible reasons include that this arc has moved away from the line of sight to the pulsar, or it may be due to significant noise interference. Further research on this issue will require longer continuous observations to obtain data with higher signal-to-noise ratios. The scintillation arc we observed was relatively diffuse (as shown in Figure 4), resulting in larger errors in the curvature measurements. The limited dataset of 4 observations reduces the precision of our parameter-fitting results. We hope to conduct more observations of pulsar B1842+14 in the future to improve the modeling accuracy. We modeled the annual modulation of the scintillation arc (as shown in Figure 10). Through modeling the annual modulation, we determined that the fractional distance $s$ of the scintillation screen is $0.85^{+0.05}_{-0.10}$, corresponding to a distance of $0.24^{+0.21}_{-0.12}$ kpc; the scattering direction $\zeta$ is $122^{+11}_{-12}$ ; and the velocity component along the scattering direction $v_{\text{scr}}$ is $19^{+11}_{-11}$ km/s. Yao et al. (2020) reported a result of 0.3 $\pm$ 0.2 kpc for this arc based on a single observation assuming $\mathbf{n}_{\text{screen}}$ aligns with $\mathbf{v}_{\text{eff}}$ and $v_{\text{screen}}$ = 0. The large uncertainty for these screen distance measurements is mainly attributed to the uncertainty in the pulsar distance. Accurate pulsar distance measurements are vital for obtaining more precise screen distance determinations.
IV.3 PSR B2021+51
@@
We discovered at least 6 scintillation arcs in the observation data of pulsar B2021+51 (see Table 5), with 6 scintillation arcs identifiable in the observation data from September 17, 2022, as shown in Figure 5. Previously, Stinebring et al. (2022) observed 2 scintillation arcs using the Green Bank Telescope.
@@
Among these, the scintillation arc E was the brightest. Our model fitting yielded the following three parameters for this scintillation screen: a fractional distance $s$ of $0.556^{+0.041}_{-0.038}$ (corresponding to a distance of $887^{+167}_{-132}$ pc), a scattering direction of $20.3^{+4.3}_{-4.3}$ , and a projected velocity of $12.1^{+3.7}_{-3.8}$ km/s. For scintillation arc D, we also successfully fitted all three parameters of the scintillation screen.
@@
IV.4 Distribution and potential sources of scintillation screens
@@
Only three potential screens (alternative solutions for B, C, and E of PSR B1237+25) might reside within the Local Bubble, and these represent just one of two possible solutions for each screen, with the alternative placements being at much greater distances. These extremely nearby solutions, if physical, would be associated with very local interstellar structures at distances of only << 1 pc. Scintillation screens close to Earth usually produce arcs with large annual curvature variation, depending on the pulsar velocity. When the annual curvature variation is large, it makes it relatively straightforward to resolve the distance degeneracy through observations at multiple epochs throughout the year.
Screens E and D for PSR B2021+51 are located far from both the Local Bubble and the pulsar itself, and thus appear unrelated to the Local Bubble or pulsar local environments, suggesting they originate from general ISM structures distributed along the line of sight.
PSR B1237+25 and PSR B2021+51 have close screens that may be associated with their local environments. However, the spin-down luminosities $\dot{E}$ of PSR B1237+25 and PSR B2021+51 are low (1.4 $\mathrm{\times 10^{31}ergs^{-1}}$ and 8.2 $\mathrm{\times 10^{32}ergs^{-1}}$, respectively), so they are unlikely to have observable bow shocks. If bow shocks were present, the bow shock radius along the line of sight would be $\lesssim$ O(1000) AU for PSR B1237+25, and $\lesssim$ O(10000) AU for PSR B2021+51 (more details in APPENDIX B). However, we only know that the closest screen to PSR B1237+25 is at a distance $\lesssim$ 40 pc (see Table 6), while that for PSR B2021+51 is $\lesssim$ 70 pc (see Table 6), so we cannot determine whether the screens are at the bow shocks. If the closest screens are at their respective bow shocks, the relative velocity between the pulsars and the screens in the main scattering direction ($v_{\mathrm{psr},\zeta}-v_{\mathrm{screen},\zeta}$) would be O(1) $\mathrm{km~s^{-1}}$ (APPENDIX B).
V Conclusion
@@
Our findings reveal that these scintillation phenomena occur throughout various regions of interstellar space between Earth and the observed pulsars. This distribution pattern indicates that electron density fluctuations capable of causing such scintillation are widespread throughout interstellar space. Several screens in our sample appear at distances that coincide with the Local Bubble boundary, particularly the brightest scintillation arc in the direction of PSR B1237+25. When confirmed to be associated with the Local Bubble boundary, the distance measurement of such scintillation screens would provide a valuable constraint on the Local Bubble shape at high Galactic latitudes. For screens located in close proximity to PSR B1237+25 and PSR B2021+51, the distance uncertainties and the uncertainty about the presence of bow shocks prevent us from determining whether these screens are associated with the pulsars’ local environments.
@@
VI Acknowledgments
This work made use of the data from FAST (Five-hundred-meter Aperture Spherical radio Telescope). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences. We are grateful to the anonymous referee for careful and insightful comments that improved this work. Xun Shi acknowledges helpful discussions with Kejia Lee and Yulan Liu, and thanks James McKee for discussions regarding observation targets. This work is supported by the National Natural Science Foundation of China grant No. 12373025. Yuanshang Huang is supported by Scientific Research and Innovation Project of Postgraduate Students in the Academic Degree of Yunnan University and the Project of Yunnan Provincial Department of Education Science Research Fund (Grant No. KC-252513932). Jumei Yao was supported by the National Science Foundation of Xinjiang Uygur Autonomous Region (2022D01D85), the Major Science and Technology Program of Xinjiang Uygur Autonomous Region (2022A03013-2), the Tianchi Talent Project, the CAS Project for Young Scientists in Basic Research (YSBR-063), the Tianshan Talents Program (2023TSYCTD0013), and the Chinese Academy of Sciences (CAS) “Light of West China” Program (No. xbzg-zdsys-202410 and No. 2022-XBQNXZ-015). Weiwei Zhu is supported by the National SKA Program of China (No. 2020SKA0120200), and the National Natural Science Foundation grant (No. 12041303). Yonghua Xu is supported by the National Natural Science Foundation (grant No. 12573104), the National Key R&D Program of China (grant No. 2022YFC2205203), and the CAS “Light of West China” Program.
@@ VI Acknowledgments
Appendix A Testing for arc apex offsets
Two mechanisms can cause the apex of a parabolic arc to be offset from the origin: (1) the scintillation screen has an electron density gradient transverse to the line of sight, or (2) the scattered images do not lie along the line of sight. However, the arcs we analyze show no measurable apex offset from the origin, indicating that the screens in our study do not exhibit significant density gradients and that their morphologies are consistent with the “parallel stripes model” (Shi, 2021). To illustrate that the scintillation arc apexes in our study show no significant offset from the origin, we consider the examples shown in Figures 2, 3, 4, and 5 (corresponding to observations of PSR B1237+25 on 2020 January 17, PSR B1237+25 on 2023 July 18, PSR B1842+14 on 2022 October 20, and PSR B2021+51 on 2021 September 17, respectively). The curvature measurement uncertainties for the main arcs in these four figures (the G arc for PSR B1237+25 and the E arc for PSR B2021+51) are 3.0, 3.7, 92, and 78 $\mathrm{m}^{-1}\,\mathrm{mHz}^{-2}$, respectively (see Table 3, 4, and 5). We performed tests without reflection padding; instead, a Hanning window was applied to the outer 10% of each edge before computing the FFT, while accounting for arc asymmetry in the fits. The differences between the most probable curvature values obtained from this alternative approach and those from reflection padding are 0.3, 0.5, 12, and 16 $\mathrm{m}^{-1}\,\mathrm{mHz}^{-2}$, respectively. These differences are smaller than the curvature measurement uncertainties.
Appendix B Estimating bow shock radii
We calculate the bow shock stand-off radius using the formula provided by Ocker et al. (2024) (based on Wilkin 1996 and Chatterjee and Cordes 2002):
$\begin{aligned} R_{0}&=\sqrt{\frac{\dot{E}}{4\pi c\rho v_{*}^{2}}}\\ &\approx 225\,\mathrm{au}\times\left[\left(\frac{\dot{E}}{10^{33}\,\mathrm{erg\,s}^{-1}}\right)\right.\\ &\quad\quad\quad\left.\times\left(\frac{n_{H}}{\mathrm{cm}^{-3}}\right)^{-1}\left(\frac{v_{*}}{100\,\mathrm{km\,s}^{-1}}\right)^{-2}\right]^{1/2},\end{aligned}$ (B1)
where $\dot{E}$ is the pulsar spin-down luminosity, $c$ is the speed of light, $\rho$ is the ISM density, $v_{*}$ is the pulsar velocity, and $n_{H}$ is the ISM hydrogen number density. $\dot{E}$ is given by $\dot{E}=4\pi^{2}I\dot{P}/P^{3}$, where $I\equiv 10^{45}\,\mathrm{g\,cm}^{2}$ is the pulsar moment of inertia. The spin-down luminosities of PSR B1237+25 and PSR B2021+51 are $1.4\times 10^{31}\,\mathrm{erg\,s}^{-1}$ and $8.2\times 10^{32}\,\mathrm{erg\,s}^{-1}$, so they are unlikely to have observable bow shocks. Assuming bow shocks are present, we use the formula from Chatterjee and Cordes (2002) describing the shape of a thin-shell bow shock to calculate the assumed bow shock radius along the line of sight:
$\begin{aligned} R(\theta)=R_{0}\csc\theta\sqrt{3(1-\theta\cot\theta)}.\end{aligned}$ (B2)
This equation is in polar coordinates, where $R(\theta)$ is the bow shock radius, and $\theta$ is the polar angle from the axis along the pulsar velocity direction. We assume that the angle between the pulsar velocity direction and the sky plane $\alpha$ is between -50° and 50°, in which case the pulsar velocity is $v_{*}=v_{\perp}/\cos\alpha$ (where $v_{\perp}$ is the pulsar’s transverse velocity), and $\theta$ ranges from 40° to 140°. We also assume that the ISM hydrogen number density $n_{H}$ ranges from $0.001$ to $0.1\,\mathrm{cm}^{-3}$. Using the above assumptions, we obtain from Equations B1 and B2 that the bow shock radius along the line of sight would be between 19 AU and 912 AU for PSR B1237+25, and 571 AU to 27563 AU for PSR B2021+51. However, we can only know that the closest screen to PSR B1237+25 is at a distance no greater than $\sim$ 40 pc (see Table 6), while that for PSR B2021+51 is no greater than $\sim$ 70 pc (see Table 6), so we cannot determine whether the screens are at the bow shocks. Although our estimated range for the bow shock radius is very large due to the substantial uncertainties in the assumptions, it still tells us that the bow shock radius is much smaller than the uncertainty in the distance from the screen to the pulsar. If the screen is very close to the pulsar ($s\ll 1$), the curvature of the scintillation arc can be approximated from Equations 3 and 10 as:
$\eta\approx\frac{d_{\text{screen-psr}}}{2({v_{\mathrm{psr},\zeta}-v_{\mathrm{screen},\zeta}})^{2}},$ (B3)
where $d_{\text{screen-psr}}$ is the distance between the screen and the pulsar, and $v_{\mathrm{psr},\zeta}$ and $v_{\mathrm{screen},\zeta}$ are the projections of the pulsar velocity and screen velocity onto the main scattering direction of the screen, respectively. If the closest screen to PSR B1237+25 is at the bow shock, we obtain from Equation B3 that $v_{\mathrm{psr},\zeta}-v_{\mathrm{screen},\zeta}$ is no greater than $5\,\mathrm{km\,s}^{-1}$, while for the closest screen to PSR B2021+51 it is no greater than $4\,\mathrm{km\,s}^{-1}$.
Appendix C Supplementary figures
Figure 8 illustrates the effectiveness of the second masking and reflection padding methods. Figures 9, 10, and 11 show the MCMC fitting results for the brightest arcs of pulsars B1237+25, B1842+41, and B2021+51, respectively. Figures 12, 13, and 14 present the posterior distributions of the scintillation screen parameters that produce the brightest arcs of pulsars B1237+25, B1842+41, and B2021+51, respectively, as well as the posterior distributions of the pulsar parallax and proper motion.
Figure 8: Secondary spectra of PSR B1237+25 on January 17, 2020, showing the effects of different preprocessing methods applied to the dynamic spectrum. Panel A: secondary spectrum obtained from the dynamic spectrum with both second masking and reflection padding applied. Panel B: only reflection padding applied. Panel C: only second masking applied. Panel D: neither second masking nor reflection padding applied. For the dynamic spectra without reflection padding, a Hanning window was applied to the outer 10% of each edge. We compare the background noise levels of each panel using the mean power away from the arc regions minus the 99.9th percentile power value: Panel A (second masking + reflection padding): -40.2 dB; Panel B (reflection padding only): -36.6 dB; Panel C (second masking only): -38.0 dB; Panel D (neither applied): -33.7 dB. Both second masking and reflection padding reduce the background noise in the resulting secondary spectrum.
@@ Appendix C Supplementary figures
Figure 12: The posterior distributions of parameters for screen G in the direction of PSR B1237+25, as well as the pulsar parallax and proper motion. In these figures, blue solid lines indicate the median, vertical dashed lines indicate 1 $\sigma$ confidence levels, and contours show 1 $\sigma$, 2 $\sigma$, and 3 $\sigma$ confidence levels.
Figure 13: The posterior distributions of parameters for the screen in the direction of PSR B1842+14, as well as the pulsar parallax and proper motion. In these figures, blue solid lines indicate the median, vertical dashed lines indicate 1 $\sigma$ confidence levels, and contours show 1 $\sigma$, 2 $\sigma$, and 3 $\sigma$ confidence levels.
Figure 14: The posterior distributions of parameters for screen E in the direction of PSR B2021+51, as well as the pulsar parallax and proper motion. In these figures, blue solid lines indicate the median, vertical dashed lines indicate 1 $\sigma$ confidence levels, and contours show 1 $\sigma$, 2 $\sigma$, and 3 $\sigma$ confidence levels.